Following data pertains to Castle State Bank. Balance Sheet: Cash 95.499, Securities 593.776, Net loans 1191.255, Net premises 297.786; Deposits 1109.808, Non-deposit borrowings 800.142, Common stock $10.452, Surplus $127.946; Income Statement: Interest Income 398.068, Interest expense 192.289, Noninterest income 51.404, Noninterest expense 95.709, Provision for loan losses 57.858, Securities gains (losses if negative) 11.275, Taxes 48.8. What is Castle State Bank's ROA? Decimalize your answers, round to four decimals. E.g., if your answer is 14.445%, record it as 0.1445. 

Note: Undivided profits is the only item missing above; it is the item that would balance the balance sheet.
Answer #1

Total Assets = Cash + Securities + Net Loans + Net Premises

= 95.499 + 593.776 + 1191.255 + 297.786 = 2,178.316

Net Income = Interest Income - Interest Expense + Non-interest income - Non-interest expense - Provision for loan losses + Securities gains - Taxes

= 398.068 - 192.289 + 51.404 - 95.709 - 57.858 + 11.275 - 48.8 = 66.091

ROA = Net Income / Total Assets = 66.091 / 2,178.316 = 0.0303, or 3.03%